A person has blue brown r green eyes .The probabilty of blue eyes is0.4The probabilty of brown eyez is 0.5 The probabilty of green eyez is 0.1.Two people chosen at random find probabilty
1)have blue eyez
2)have different coloured eyez
1
Expert's answer
2020-06-22T18:08:30-0400

1) P=P(blue)*P(blue)=0.4*0.4=0.16

2)P=1-P(blue)*P(blue)-P(brown)*P(brown)-P(green)*P(green)=1-0.16-0.25-0.01=0.58